# Asymmetric Errors

### v2

@@ 2.1.2 Comparison of the models
Dimidiated Gaussian OPAT with 2 straight lines ±$\pm 1$ Special case Discontinuity which can give problems when fitting
Distorted Gaussian OPAT with a parabola ±$\pm 0.57382$ Yes Limited range
Railway Gaussian OPAT with parabola morphing to straight lines ±$\pm 0.60467$ Yes Arbitrary smoothing
Double cubic Gaussian OPAT with two cubics morphing to straight lines ±$\pm 0.74538$ Yes Arbitrary smoothing
Symmetric beta Gaussian OPAT with polynomial morphing to straight lines ±$\pm 1$ Yes Two arbitrary tuning parameters
QVW Gaussian Gaussian with $\sigma$ linear function of cumulant ±$\pm 0.68269$ No Messy numerically
Fechner distribution Two half Gaussians of same height ±$\pm 0.21564$ No Little motivation
@@
Skew Normal Azzalini’s form ±$\pm 0.21564$ No
Maximum entropy Johnson Johnson $S_{U}$$\pm 0.34719$ No Kurtosis from maximum entropy
Log normal Takes exp of Gaussian distributed variable ±$\pm 1$ No Limited range
@@
The upper row of Figure 6 shows several models, with parameters chosen to match quantile values (on the left) and moments (on the right). For these moderate asymmetries the behaviour appears sensible and there is generally little difference between the shapes, except for the dimidiated Gaussian near the centre (where it probably doesn’t matter very much).
Figure 6: Pdfs for the different models. Panel A is for ${5.0}^{+1.1}_{-0.9}$, B is for moments: $\mu=5.0,V=2.0,\gamma=-1.0$, C is for ${5.0}^{+1.3}_{-0.7}$ and panel D for $\mu=5,V=2.0,\gamma=-3.0$.
Two examples of larger asymmetry are shown in the lower row, and differences appear. Azzalini’s skew normal distribution cannot handle these values. The Edgeworth form goes negative and develops a second bump, which is surely not physical. The distorted Gaussian shows a Jacobian peak which, again, is probably an undesired artefact. The railway Gaussian resembles the distorted but avoids the peak.
@@
As a further exploration, we can use some examples of the third original category of Section 1.1, where a Gaussian-distributed variable is transformed to make it non-Gaussian. We apply the various models to the moments of this distribution, and compare the model with the true original. Examples are shown in Figure 7, taking the square of a Gaussian-distributed variable, the square root, the exponential and the logarithm. (Different means and standard deviations were used, for presentational reasons; square roots and logarithms of negative numbers are ignored.) For small to moderate asymmetries there is general agreement among the models (apart from the dimidiated model near the centre, as before) and they match the original. The fourth example shown, D, considers the logarithm of a Gaussian variable, and is more interesting in that the models generally disagree with each other and the original.
Figure 7: Pdfs from transformations of Gaussian distributions, and the model fits to them. Panel A is the distribution for $x^{2}$, where $x$ is distributed according to $\phi(10.0,2.0)$. B is for $\sqrt{x}$ with $x$ from $\phi(10.0,3.0)$. C is $e^{x}$ with $x$ from $\phi(5.0,0.25)$ and panel $D$ is $\ln x$ with $x$ from $\phi(5.0,2.0)$.
It is strongly recommended that any calculations with asymmetric errors use (at least) two models, as a check for robustness. The dimidiated and distorted models have the advantage of simplicity, even though some aspects are clearly unrealistic. The railway model is similar to the distorted and avoids the Jacobian peak. On the other hand the Edgeworth model utilizing only three leading cumulants has an inbuilt problem in that it always gives a negative (unphysical) probability for some arguments. Azzalini’s skew normal is restricted to small asymmetries. It seems to us that the Edgeworth and skew normal models are not suitable for general-purpose use, which have to handle a wide range of possible asymmetries: they can be used if there is a specific reason for favouring them, and it is known a priori that the asymmetries will not be large.
@@ 2.2 Combination of pdf errors
The combination of two (or more) errors requires the convolution of the relevant pdfs: given $p_{x}(x)$ and $p_{y}(y)$ with $u=u(x,y)$ we require $p_{u}(u)$. In the linear approximation we assume, without loss of generality, that the values are shifted to have a nominal value of 0, $u(x,y)=u(x_{0},y_{0})+\tilde{u}(x-x_{0},y-y_{0})$ and then scaled to absorb the differentials, $x\to x_{0}+\left({\partial u\over\partial x}\right)\tilde{x}$ and $y\to y_{0}+\left({\partial u\over\partial y}\right)\tilde{y}$. Then (dropping the tildes) we have $u=x+y$ and the pdf for $u$ is the convolution of the two , as given in Equation 3.In some cases this can be done algebraically, in others it can be done numerically.
This presents a problem, in that, although the convolution of two Gaussians gives another Gaussian, this is not in general true for other functions. The convolution of two pdfs, both belonging to one of the parametrisations in Table 2, does not give a function described by that parameterisation. See Appendix A.1.1 for a detailed discussion using the dimidiated Gaussian.
     To proceed, we note that the moments $\mu,V,\gamma$, add under convolution. To provide a consistent procedure for ‘adding errors’ we can evaluate the individual moments using Equations (17), (25), or their equivalents for other models, sum them to obtain the totals, and then use Equations (19), (26), or their equivalents to find the model parameters and, if desired, the quantile parameters, that give a pdf that has these moments.
@@
An important point to note is that in all cases the nominal value shifts: this is the number shown as $\Delta$. If two asymmetric pdfs are convolved, then the median of the result is not the sum of the individual medians. A procedure for combining errors using asymmetric distributions must, to be consistent, include a shift in the central value. If practitioners are not prepared to make such an adjustment then they should perhaps reconsider the use of asymmetric errors and revert to the usual symmetric form, averaging their ${\sigma^{+}}$ and ${\sigma^{-}}$ values. The Values of ${\sigma^{+}}$ and ${\sigma^{-}}$ in the table, and returned by the software, give, within the framework of the model concerned, the 68% central region around the nominal point: when the $\Delta$ shift is applied they should be adjusted accordingly.
@@ 2.3 Why the ‘usual procedure’ is wrong
It is common practice (though we have not been able to find a reference) to combine asymmetric errors by combining all the positive errors in quadrature, and likewise all the negative errors, and then using the dimidiated (“bifurcated”) Gaussian. This is obviously wrong. Suppose one has $N$ sources of error, and that they all have the same positive error ${\sigma^{+}}$ and negative error ${\sigma^{-}}$. The combined error given by this procedure will have positive error $\sqrt{N}{\sigma^{+}}$ and negative error $\sqrt{N}{\sigma^{-}}$. The pdf of the combination has a width that increases like $\sqrt{N}$ but does not change its shape. But this contradicts the Central Limit Theorem, which declares that at large $N$ all distributions tend towards a Gaussian (symmetric) shape.
The flaw in the logic is as follows: when two random variables, described (for the sake of the example) by dimidiated Gaussians, are combined, there is a 25% chance that they will both fluctuate upwards. That is indeed described, within the framework of this model, by a 2-D Gaussian for which the standard deviation of the sum is the sum in quadrature of the two positive $\sigma$. Likewise there is a 25% chance that they will both go negative, and that is described by $\sqrt{{\sigma^{-}_{1}}^{2}+{\sigma^{-}_{2}}^{2}}$. But there is a 50% chance that one will go up and the other will go down. This is neglected by the ‘usual procedure’, and it is this filling in of central values that lets the Central Limit Theorem do its work.
@@ 2.4 Combination of pdf results
As discussed in Section 1.1, the use of likelihoods to combine results is a very standard procedure and their use to produce a combined error is less mainstream, whereas for pdfs the case is reversed. In fact, for pdfs the combination of results can be considered as an instance of the combination of errors. The measurements are all of the same quantity and combined by taking a weighted meanrather than distinct quantities combined with some general function.
@@
Continuing with this example, we use a toy Monte Carlo to generate pairs of ‘samples’ in $x$ according to a random Gaussian process, combine the two, and histogram the results. This is shown in Figure 8. The mean, variance and skewness can be found from each histogram and compared with the values expected from the quoted results, shown in Table 5. It can be seen that there is good agreement for the variances in all cases and fair agreement for the skewness, except that the dimidiated Gaussian underestimates the skewness of the pdf of the combination, because the dimidiated model is less influenced by extreme values than the others.
Figure 8: Combination of results using pdfs.
@@
For a given measurement ${x_{i}}^{+{\sigma^{+}_{i}}}_{-{\sigma^{-}_{i}}}$ we can express its compatibility with some quoted $x_{c}$ using the p-value or its equivalent expression in terms of sigmas. With the familiar Gaussian, we would say that a measurement of $12.7\pm 0.1$ was 5 sigma away from some nominal value of $12.2$, and accordingly treat it with strong reservation. A similar approach for asymmetric errors will use ${\sigma^{+}_{i}}$ and ${\sigma^{-}_{i}}$, but care must be taken over the direction of the deviation. Suppose the proposed true value is $12.2$ and the measurement is ${12.7}^{+0.1}_{-0.2}$. Using the dimidiated model, this deviation is a 5 sigma, not 2.5 sigma, as the quoted measurement implies that a process which may go up a little or down a lot has given a result of 12.7. For other models the arithmetic is not so simple, but the p-value can readily be determined from the pdf.
@@ 3.1 Modelling non-parabolic likelihoods
Two models which are simple to use and turn out to be generally applicable arise from the suggestion by Bartlett that a likelihood function is described by a Gaussian whose width varies as a function of the argument $a$ [16, 17, 13]. With such an assumption one might suppose that the variation is linear, either for the standard deviation (‘linear sigma model’) or the variance (‘linear variance model’).
@@
If any information is provided, or can be deduced, about the form of the likelihood then that should be used to guide the choice of model – there is no general ‘best model’, but some models do better for particular classes of result. By the same token, if a result with an asymmetric error is quoted any further available information should be provided. If all the information available to a meta-analyst is the 3 numbers of the reported measurement, then the wide range of available models can potentially give a wide range of outcomes.
@@ 3.1.1 Comparison of models and recommendations
As an illustration we take the result $a={5.0000}^{+2.5811}_{-1.9159}$, which would be reported from a Poisson measurement of 5 events. This is shown in Figure 10. The true likelihood is $L(a)=e^{-a}{a^{5}\over 5!}$, and we compare this with the results of the likelihoods given by the models. All curves, by construction, peak at $a=5$ and go through the $\Delta\ln L=\pm{1\over 2}$ points, and their interpolations in the central region are very similar. At larger values their behaviours are very different. The cubic turns over, which is unacceptable. The quartic provides a better match to the original Poisson likelihood, especially on the negative side. The logarithmic form does fairly well, the Poisson does perfectly, as it should in this particular example. The two Bartlett forms, labelled Linear Variance and Linear Sigma,both do well, the linear variance is somewhat better than the linear sigma. The PDG form does badly, as it assumes a Gaussian behaviour outside the central interval. The Edgeworth form does badly at both small and large values, and is troubled by logarithms of negative numbers at large values and/or asymmetries. The skew normal is not as bad as Edgeworth’s approximation, but it does not do very well. These comparisons assume the fundamental $a={5.0000}^{+2.5811}_{-1.9159}$ measurement comes from a Poisson measurement: if it does not, then conclusions might be different.
@@
On the basis of these two examples it looks as if the two Bartlett forms generally behave well. There is not a lot to choose between them, but one will not go far wrong using the linear variance with the linear sigma as a sanity check, though the other way round would be quite acceptable – particularly if the measurement is basically a Poisson process. Other models may be useful in particular applications.
@@ 3.2 Combination of results using likelihood
If one has several results measuring the same quantity then the total likelihood is the product of the individual likelihoods, and the log likelihood is just the sum of the individual log likelihoodsFrom that one can read off the overall $\hat{a}$ and the $\Delta\ln L=-{1\over 2}$ errors. If we do not know the full forms for the likelihoods, but are only given the peak values and errors, then we can use the same approach, but modelling the likelihoods by one of the methods above. In general a numerical maximisation can be used to give a solution. For the two Bartlett models this can be done efficiently: the likelihood is
$\ln L(a)=-{1\over 2}\sum_{i}\left({a-\hat{a}_{i}\over\sigma_{i}(a)}\right)^{2}$ (10)
@@
The equivalent for the linear v‘ariance model is
@@
Table 7: Results of combining the same 3 results according to different models.
@@
It is worth pointing out how the combined error depends on the values being combined, as well as their quoted errors. If one of the values in the combination is changed this affects not only the combined central value but also the quoted error. In the symmetric case the quoted errors are independent of the values.

### v3

@@ 2.1.2 Comparison of the models
Dimidiated Gaussian OPAT with 2 straight lines ±1$\pm$ 1± 1 Special case Discontinuity which can give problems when fitting
Distorted Gaussian OPAT with a parabola ±0.57382$\pm$ 0.57382± 0.57382 Yes Limited range
Railway Gaussian OPAT with parabola morphing to straight lines ±0.60467$\pm$ 0.60467± 0.60467 Yes Arbitrary smoothing
Double cubic Gaussian OPAT with two cubics morphing to straight lines ±0.74538$\pm$ 0.74538± 0.74538 Yes Arbitrary smoothing
[…]
QVW Gaussian Gaussian with $\sigma$ linear function of cumulant ±0.68269$\pm$ 0.68269± 0.68269 No Messy numerically
Fechner distribution Two half Gaussians of same height ±0.21564$\pm$ 0.21564± 0.21564 No Little motivation
@@
Skew Normal Azzalini’s form ±0.21564$\pm$ 0.21564± 0.21564 No
Maximum entropy Johnson Johnson $S_{U}$0.34719$\pm$ 0.34719± 0.34719 No Kurtosis from maximum entropy
Log normal Takes exp of Gaussian distributed variable ±1$\pm$ 1± 1 No Limited range
@@
The upper row of Figure 6 shows several models, with parameters chosen to match quantile values (on the left) and moments (on the right). For these moderate asymmetries the behaviour appears sensible and there is generally little difference between the shapes and the curves coincide,except for the dimidiated Gaussian near the centre (where it probably doesn’t matter very much).
Figure 6: Pdfs for a fewdifferent models, listed in the legends.Panel A is for ${5.0}^{+1.1}_{-0.9}$, B is for moments: $\mu=5.0,V=2.0,\gamma=-1.0$, C is for ${5.0}^{+1.3}_{-0.7}$ and panel D for $\mu=5,V=2.0,\gamma=-3.0$.
Two examples of larger asymmetry are shown in the lower row, and differences appear. Azzalini’s skew normal distribution cannot handle these values. The Edgeworth form goes negative and develops a second bump, which is surely not physical. The distorted Gaussian shows a Jacobian spike which, again, is probably an undesired artifact. The railway Gaussian resembles the distorted but avoids the spike.
@@
As a further exploration, we can use some examples of the third original category of Section 1.1, where a Gaussian-distributed variable is transformed to make it non-Gaussian. We apply the various models to the moments of this distribution, and compare the model with the true original. Examples are shown in Figure 7, taking the square of a Gaussian-distributed variable, the square root, the exponential and the logarithm. (Different means and standard deviations were used, for presentational reasons; square roots and logarithms of negative numbers are ignored.) For small to moderate asymmetries there is general agreement among the models (apart from the dimidiated model near the centre, as before) and they match the original. The fourth example shown, panel D, considers the logarithm of a Gaussian variable, and is more interesting in that the models generally disagree with each other and the original.
Figure 7: Pdfs from transformations of Gaussian distributions, and the model fits to them. Panel A is the distribution for $x^{2}$, where $x$ is distributed according to $N(10.0,2.0)$. B is for $\sqrt{x}$ with $x$ from $N(10.0,3.0)$. C is $e^{x}$ with $x$ from $N(5.0,0.25)$ and panel $D$ is $\ln x$ with $x$ from $N(5.0,2.0)$.
It is strongly recommended that any calculations with asymmetric errors use (at least) two models, as a check for robustness. The dimidiated and distorted models have the advantage of simplicity, even though some aspects are clearly unrealistic. The railway model is similar to the distorted and avoids the Jacobian spike. On the other hand the Edgeworth model utilizing only three leading cumulants has an inbuilt problem in that it always gives a negative (unphysical) probability for some arguments. Azzalini’s skew normal is restricted to small asymmetries. It seems to us that the Edgeworth and skew normal models are not suitable for general-purpose use, which have to handle a wide range of possible asymmetries: they can be used if there is a specific reason for favouring them, and it is known a priori that the asymmetries will not be large.
@@ 2.2 Combination of pdf errors
The combination of two (or more) errors requires the convolution of the relevant pdfs: given $p_{x}(x)$ and $p_{y}(y)$ with $u=u(x,y)$ we require $p_{u}(u)$. One assumes that the errors are ‘small’ so the linear approximation can be used. Near the nominal point $(x_{0},y_{0})$ we write $u(x,y)\approx u_{0}+\left({\partial u\over\partial x}\right)(x-x_{0})+\left({\partial u\over\partial y}\right)(y-y_{0})$. Scaling to absorb the differentials we write $\tilde{u}=u-u_{0},\tilde{x}=\left({\partial u\over\partial x}\right)(x-x_{0}),\tilde{y}=\left({\partial u\over\partial y}\right)(y-y_{0})$, which gives $\tilde{u}=\tilde{x}+\tilde{y}$. So within the limits of the approximation we can express any general $u=u(x,y)$ as a simple sum and for simplicity in the following algebra for error manipulations we will work with $u=x+y$, knowing that the general form can be reclaimed when and if desired. Using this, $p_{u}(u)$ is given by Equation 3.In some cases this can be evaluated algebraically, in others it needs to be done numerically.
This presents a problem, in that, although the convolution of two Gaussians gives another Gaussian, this is not in general true for other functions. The convolution of two pdfs, both belonging to one of the parameterisations in Table 2, does not give a function described by that parameterisation. See Appendix A.1.1 for a detailed discussion using the dimidiated Gaussian.
      To proceed, we note that the moments $\mu,V,\gamma$ add under convolution. To provide a consistent procedure for ‘adding errors’ we can evaluate the individual moments using Equations (17), (25), or their equivalents for other models, sum them to obtain the totals, and then use Equations (19), (26), or their equivalents, to find the model parameters and, if desired, the quantile parameters that give a pdf that has these moments.
@@
An important point to note is that in all cases the nominal value shifts: this is the number shown as $\Delta$. If two asymmetric pdfs are convolved, then the median of the result is not the sum of the individual medians. A procedure for combining errors using asymmetric distributions must, to be consistent, include a shift in the central value. If practitioners are not prepared to make such an adjustment, then they should perhaps reconsider the use of asymmetric errors and revert to the usual symmetric form, averaging their ${\sigma^{+}}$ and ${\sigma^{-}}$ values. The values of ${\sigma^{+}}$ and ${\sigma^{-}}$ in the table, and returned by the software, give, within the framework of the model concerned, the 68% central region around the nominal point: when the $\Delta$ shift is applied the region boundaries should be adjusted accordingly.
@@ 2.3 Why the ‘usual procedure’ is wrong
It is common practice (though we have not been able to find a reference) to combine asymmetric errors by combining all the positive errors in quadrature, and likewise all the negative errors, and then using the dimidiated (“bifurcated”) Gaussian. This is obviously wrong. Suppose one has $N$ sources of error, and that they all have the same positive error ${\sigma^{+}}$ and negative error ${\sigma^{-}}$. The combined error given by this procedure will have positive error $\sqrt{N}{\sigma^{+}}$ and negative error $\sqrt{N}{\sigma^{-}}$. The pdf of the combination has a width that increases like $\sqrt{N}$ but does not change its shape. But this contradicts the Central Limit Theorem, which declares that at large $N$ the distribution of a sum tends to a Gaussian (symmetric) shape, whatever the distributions of the components.
@@ 2.4 Combination of pdf results
As discussed in Section 1.1, the use of likelihoods to combine results is a very standard procedure and their use to produce a combined error is less mainstream, whereas for pdfs the case is reversed. In fact, for pdfs the combination of results can be considered as an instance of the combination of errors. The measurements are all of the same quantity and combined by taking a weighted mean, rather than distinct quantities combined with some general function.
@@
Continuing with this example, we use a toy Monte Carlo to generate pairs of ‘samples’ in $x$ according to a random Gaussian process as described above,combine the two, and histogram the results. This is shown in Figure 8. The mean, variance and skewness can be found from each histogram and compared with the values expected from the quoted results, shown in Table 5. It can be seen that there is good agreement for the variances in all cases and fair agreement for the skewness, except that the dimidiated Gaussian underestimates the skewness of the pdf of the combination, because the dimidiated model is less influenced by extreme values than the others.
Figure 8: Example, using Monte Carlo simulations with 10,000 samples, of combination of results using pdfs, taking two random Gaussian numbers and squaring them, both being ascribed errors ${}^{+7.57}_{-6.57}$, and then combined using different pdf models. From such distributions the quantities $V_{MC}$ and $\gamma_{MC}$ in Table 5 are calculated.
@@
For a given measurement ${x_{i}}^{+{\sigma^{+}_{i}}}_{-{\sigma^{-}_{i}}}$ we can express its compatibility with some quoted $x_{c}$ using the p-value or its equivalent expression in terms of sigmas. With the familiar Gaussian, we would say that a measurement of 12.7±0.112.7$\pm$ 0.112.7 ± 0.1 was 5 sigma away from some nominal value of 12.212.212.2, and accordingly treat it with strong reservation. A similar approach for asymmetric errors will use ${\sigma^{+}_{i}}$ and ${\sigma^{-}_{i}}$, but care must be taken over the direction of the deviation. Suppose the proposed true value is 12.212.212.2 and the measurement is ${12.7}^{+0.1}_{-0.2}$. Using the dimidiated model, this deviation is a 5 sigma, not 2.5 sigma, as the quoted measurement implies that a process which may go up a little or down a lot has given a result of 12.7. For other models the arithmetic is not so simple, but the p-value can readily be determined from the pdf.
@@ 3.1 Modelling non-parabolic likelihoods
Two models which are simple to use and turn out to be generally applicable arise from the suggestion by Bartlett that a likelihood function is described by a Gaussian whose width varies as a function of the argument $a$ [16, 17, 13]. With such an assumption one might suppose that the variation is linear, either for the standard deviation (‘linear sigma model’) or the variance (‘linear variance model’)
@@
If any information is provided, or can be deduced, about the form of the likelihood then that should be used to guide the choice of model — there is no general ‘best model’, but some models do better for particular classes of result. By the same token, if a result with an asymmetric error is quoted, any further available information should be provided. If all the information available to a meta-analyst is the 3 numbers of the reported measurement, then the wide range of available models can potentially give a wide range of outcomes.
@@ 3.1.1 Comparison of models and recommendations
As an illustration we take the result $a={5.0000}^{+2.5811}_{-1.9159}$, which would be reported from a Poisson measurement of 5 events. This is shown in Figure 10. The true likelihood is $L(a)=e^{-a}a^{5}/5!$, and we compare this with the results of the likelihoods given by the models. All curves, by construction, peak at $a=5$ and go through the $\Delta\ln L=\pm{1\over 2}$ points, and their interpolations in the central region are very similar. At larger values their behaviours are very different. The cubic turns over, which is unacceptable. The quartic provides a better match to the original Poisson likelihood, especially on the negative side. The logarithmic form does fairly well, the Poisson does perfectly, as it should in this particular example. The two Bartlett forms, labelled Linear Variance and Linear Sigma, both do well, the linear variance is somewhat better than the linear sigma. The PDG form does badly, as it assumes a Gaussian behaviour outside the central interval. The Edgeworth form does badly at both small and large values, and is troubled by logarithms of negative numbers at large values and/or asymmetries. The skew normal is not as bad as Edgeworth’s approximation, but it does not do very well. These comparisons assume the fundamental $a={5.0000}^{+2.5811}_{-1.9159}$ measurement comes from a Poisson measurement: if it does not, then conclusions might be different.
@@
On the basis of these two examples it looks as if the two Bartlett forms generally behave well. There is not a lot to choose between them, but one will not go far wrong using the linear variance with the linear sigma as a sanity check, though the other way round would be quite acceptable — particularly if the measurement is basically a Poisson process. Other models may be useful in particular applications.
@@ 3.2 Combination of results using likelihood
If one has several results measuring the same quantity, then the total likelihood is the product of the individual likelihoods, and the log likelihood is just the sum of the individual log likelihoods. From that, one can read off the overall $\hat{a}$ and the $\Delta\ln L=-{1\over 2}$ errors. If we do not know the full forms for the likelihoods, but are only given the peak values and errors, then we can use the same approach, but modelling the likelihoods by one of the methods above. In general a numerical maximisation can be used to give a solution. For the two Bartlett models this can be done efficiently: the likelihood is
$\ln L(a)=-{1\over 2}\sum_{i}\left({a-\hat{a}_{i}\over\sigma_{i}(a)}\right)^{2},$ (10)
@@
The equivalent for the linear variance model is
@@
Table 7: Results of combining the same 3 results, ${1.9}^{+0.7}_{-0.5}$, ${2.4}^{+0.6}_{-0.8}$ and ${3.1}^{+0.5}_{-0.4}$,according to different models.
@@
It is worth pointing out how the combined error depends on the values being combined, as well as their quoted errors. If one of the values in the combination is changed, this affects not only the combined central value but also the quoted error. In the symmetric case the quoted errors are independent of the values.